# Dequantization and Color Transfer with Diffusion Models

### v4

@@ Abstract
We demonstrate an image dequantizing diffusion model that enables novel image edits on natural images. We propose operating on quantized images because they offer easy abstraction for patch-based edits and palette transfer. In particular, we show that color palettes can make the output of the diffusion model easier to control and interpret. We first establish that existing image restoration methods are not sufficient, such as JPEG noise reduction models. We then demonstrate that our model can generate natural images that respect the color palette the user asked for. For palette transfer, we propose a method based on weighted bipartite matching. We then show that our model generates plausible images even after extreme palette transfers, respecting user query. Our method can optionally condition on the source texture in part or all of the image. In doing so, we overcome a common problem in existing image colorization methods that are unable to produce colors with a different luminance than the input. We evaluate several possibilities for texture conditioning and their trade-offs, including luminance, image gradients, and thresholded gradients, the latter of which performed best in maintaining texture and color control simultaneously. Our method can be usefully extended to another practical edit: recoloring patches of an image while respecting the source texture. Our procedure is supported by several qualitative and quantitative evaluations.
@@ 1 Introduction
We describe image editing procedures for modifying the color of natural images. Our editing procedures allow the artist to apply natural vector art edits to pixel art. For example, an artist can modify the color palette of an image - including the colors in the palette and which patches are assigned to each color. Spatial layout control of reference image based color transfer is not a common capability of existing methods. We use diffusion models to “snap” the artist’s changes to realistic images.
@@ 2 Related Work
Image Synthesis: There is a rich history of image synthesis with GANs [16, 19, 34] and diffusion models [5, 33]. The quality of the generated distributions is often quite good, with problems like aliasing [18] and sparse datasets [17] well-explored. Techniques have been developed to condition the generation on additional inputs like depth or normals [7, 40].
@@
We are not aware of an inpainting method that allows the user to specify the approximate color of the inpainted region as an RGB triple with potentially different luminance (previous work only conditions on a text caption e.g. “red bird,” or a limited number of colors and objects at low resolution [20]). We argue this is an important image editing activity and useful to have in practice. In this work, we show this edit can be accomplished with ControlNet, and any segmenter like SegmentAnything  [21, 40]; it’s likely other inpainting methods can synthesize images conditioned on color, although no one has shown this capability.
Palette mapping can be seen as a variant of style transfer, though we deprecate this view. Style transfer methods (e.g. [13, 22, 26, 31, 9, 10]) do change palettes, but change the texture of the image as well; in contrast we want to change only the colors. Some methods require finetuning a style transfer model for each new style [42, 8]; others can generalize to new styles [9, 10, 35, 4, 14, 11]. [39] uses text prompts rather than images to guide the style transfer. There is a small literature of palette transfer methods (with a review, [25]). Chang et al. demonstrate the value of quantizing a palette in a wholly interactive method. A number of automatic methods are framed as a warp that matches color spaces. Reinhard et al. match moments in Lab color space [30]; they obtain improved transfer by segmenting image and reference and matching segments. Wu et al use an explicit semantic representation to match (so flowers to flowers and grass to grass, say) [37]. Pitié et al. warp the color space using repeated 1D projections and a form of histogram matching [29]; Hwang et al. use moving least squares [12]. Cho et al. train a network to apply a specified palette to a natural image using color augmentations [3]. While palette mapping could be viewed as a colorization problem (decolor, then colorize with the specified palette), we are not aware of colorization algorithms that can be forced to use a restricted palette. Existing style transfer methods typically disrupt texture patterns, as well as change color in uncontrolled fashion. Basioti et al. de-quantize small face images using a GAN [1].
@@ 3 Method
Dataset: We collect 1.8M images from LAION-Aesthetic [^1] https://laion.ai/. for training, and hold-out a validation and test set. This dataset spans natural images capturing a wide range of lighting conditions, subjects and textures. An array of stylistic artwork is represented, such as digital paintings.
@@
Inference - Palette Transfer: Once we establish that our method can effectively dequantize, then we can use color palettes as an effective abstraction for image editing. For example, we can edit one or more colors in the palette, or which patches get assigned which color. Further, our method can dequantize images over a wide range of colors in the palette, supporting varying levels of abstraction. To perform a palette transfer, we can quantize a source image and extract its palette. Then we select a target color map, which could be extracted from another image. Then we perform weighted bipartite matching, assigning each source color the most similar target color, ending up with a 1:1 correspondence. Given this modified quantized image, we can then dequantize using our diffusion model to obtain the palette-transferred result. The exact palette matching procedure is up to the artist - our focus in this work is to show that our model respects the palette that was asked for, even extreme palette transfers. We show results for two color-matching algorithms. We call matching based on the most similar color as the $\mathtt{color}$ method; by negating the pairwise distance matrix between source and target colors, we can effectively match based on most dissimilar color, which results in interesting effects ($\mathtt{negative-color}$).
@@ 4 Results & Discussion
Dequantization with Texture: In Fig. 4, we evaluate our models when texture conditioning is available. (top) We are quantizing the GT image and attempting to reconstruct it from its palette, along with additional texture conditioning (i.e. thresholded gradient, gradient, luminance). (bottom) In the case of palette transfers, we do not know the GT image; thus to evaluate, we can quantize the generated image and compare it with the input (which is a quantized image). Observe how the quality improves with more colors in the palette. Thus, the Palette-SSIM metric is a reasonable proxy when GT is not available, but it’s not perfect. Notice that $\mathsf{ours-T}$ (thresholded gradient) performs better than $\mathsf{ours-L-post}$ (which conditions on GT luminance and enforces it as post-process). Palette error has curious difficulties that follow from the behavior of the quantization algorithm. If we impose the original image’s luminance, we might change the quantization substantially. In Fig. 6 we evaluate our four methods when texture accompanies the quantized image as conditioning. The row indicates the number of colors in the palette; the first column is the quantized image. The subsequent 3 columns show conditioning texture for the 3 models. $\mathsf{ours-L-post}$ and $\mathsf{ours-L}$ condition on luminance, with the former enforcing luminance as a post-process. $\mathsf{ours-G}$ conditions on image gradients and $\mathsf{ours-T}$ conditions on thresholded gradients. All models - even those with less information available - are capable of dequantizing the input quite well, though the luminance-based models tend to achieve better color reproduction.
@@
Table 1: Palette transfer error metrics for $\mathtt{color}$ method. Even though $\mathsf{ours-T}$ performed slightly worse in texture-conditioned dequantization, it excels in palette transfers because thresholded gradients decorrelate local texture from local luminance.
@@ 6 Additional Patch Editing Examples
Quantitative evaluation We select 1000 random test images at 256-res, and pick random patches in each to remove. On average, $13.4\%$ of pixels are masked out. In Table 3, we replace the removed patch with the mean color of its support, and measure how well the model can reconstruct the image. As expected, more texture information is helpful, as shown by the strong numbers for $\mathsf{ours-L}$. However, evaluating each model’s ability to support extreme color changes tells a different story. In Table 4, we repeat this experiment, but we replace each patch with a random $(R,G,B)$ triple. Here, conditioning on texture resulted in the thresholded gradient model performing best, while luminance-conditioning performed worst. This is because conditioning on luminance prevented the network from generating colors with a different luminance that what was asked for in the patch conditioning. Note that in the absence of GT, we use palette error as a proxy error metric. This metric is not perfect, as for example the models without texture conditioning performed slightly better. However, the palette error gives a strong indication as to whether the model generated an image whose color palette matches what was requested. We use 16 colors in the palette when running this error metric.
@@
ours-T 30.68 0.9336 29.3 0.921
ours-G 38.77 0.9823 30.95 0.9315
ours-L 40.12 0.9845 29.68 0.9202
ours-L-post 39.82 0.9867 38.27 0.9834
@@
ours-T 23.56 0.8241 23.87 0.8285
ours-G 22.37 0.7964 23.58 0.8208
ours-L 21.76 0.7895 24.07 0.8347
ours-L-post 20.57 0.7739 20.53 0.7731

### v5

@@ Abstract
We demonstrate an image dequantizing diffusion model that enables novel edits on natural images. We propose operating on quantized images because they offer easy abstraction for patch-based edits and palette transfer. In particular, we show that color palettes can make the output of the diffusion model easier to control and interpret. We first establish that existing image restoration methods are not sufficient, such as JPEG noise reduction models. We then demonstrate that our model can generate natural images that respect the color palette the user asked for. For palette transfer, we propose a method based on weighted bipartite matching. We then show that our model generates plausible images even after extreme palette transfers, respecting user query. Our method can optionally condition on the source texture in part or all of the image. In doing so, we overcome a common problem in existing image colorization methods that are unable to produce colors with a different luminance than the input. We evaluate several possibilities for texture conditioning and their trade-offs, including luminance, image gradients, and thresholded gradients, the latter of which performed best in maintaining texture and color control simultaneously. Our method can be usefully extended to another practical edit: recoloring patches of an image while respecting the source texture. Our procedure is supported by several qualitative and quantitative evaluations.
@@ 1 Introduction
We describe image editing procedures for modifying the color of natural images. Our editing procedures allow artists to apply natural vector art edits to pixel art. For example, an artist can modify the color palette of an image - including the colors in the palette and which patches are assigned to each color. Spatial layout control of reference image based color transfer is not a common capability of existing methods. We use diffusion models to “snap” the artist’s changes to realistic images.
@@ 2 Related Work
Image Synthesis: There is a rich history of image synthesis with GANs [16, 19, 34, 40, 36] and diffusion models [5, 33]. The quality of the generated distributions is often quite good, with problems like aliasing [18] and sparse datasets [17] well-explored. Techniques have been developed to condition the generation on additional inputs like depth, normals [7, 45], and even 3D primitives [37, 39, 38].
@@
We are not aware of an inpainting method that allows the user to specify the approximate color of the inpainted region as an RGB triple with potentially different luminance (previous work only conditions on a text caption e.g. “red bird,” or a limited number of colors and objects at low resolution [20]). We argue this is an important image editing activity and useful to have in practice. In this work, we show this edit can be accomplished with ControlNet, and a segmenter like SegmentAnything  [21, 45]; it’s likely other inpainting methods can synthesize images conditioned on color, although no one has shown this capability.
Palette mapping can be seen as a variant of style transfer, though we deprecate this view. Style transfer methods (e.g. [13, 22, 26, 31, 9, 10]) do change palettes, but change the texture of the image as well; in contrast we want to change only the colors. Some methods require finetuning a style transfer model for each new style [47, 8]; others can generalize to new styles [9, 10, 35, 4, 14, 11]. [44] uses text prompts rather than images to guide the style transfer.
There is a small literature of palette transfer methods (with a review, [25]). Chang et al. demonstrate the value of quantizing a palette in a wholly interactive method. A number of automatic methods are framed as a warp that matches color spaces. Reinhard et al. match moments in Lab color space [30]; they obtain improved transfer by segmenting image and reference and matching segments. Wu et al use an explicit semantic representation to match (so flowers to flowers and grass to grass, say) [42]. Pitié et al. warp the color space using repeated 1D projections and a form of histogram matching [29]; Hwang et al. use moving least squares [12]. Cho et al. train a network to apply a specified palette to a natural image using color augmentations [3]. While palette mapping could be viewed as a colorization problem (decolor, then colorize with the specified palette), we are not aware of colorization algorithms that can be forced to use a restricted palette. Existing style transfer methods typically disrupt texture patterns, as well as change color in uncontrolled fashion. Basioti et al. de-quantize small face images using a GAN [1].
@@ 3 Method
Dataset: We collect 1.8M images from LAION-Aesthetic [^1] https://laion.ai/. for training, and hold-out a validation and test set. This dataset spans natural images, capturing a wide range of lighting conditions, subjects, and textures. An array of stylistic artwork is represented, such as digital paintings.
@@
Inference - Palette Transfer: Once we establish that our method can effectively dequantize, then we can use color palettes as an effective abstraction for image editing. For example, we can edit one or more colors in the palette, or which patches get assigned which color. Further, our method can dequantize images over a wide range of colors in the palette, supporting varying levels of abstraction. To perform a palette transfer, we can quantize a source image and extract its palette. From there, we select a target color map, which could be extracted from another image. Then we perform weighted bipartite matching, assigning each source color the most similar target color, ending up with a 1:1 correspondence. Given this modified quantized image, we can then dequantize using our diffusion model to obtain the palette-transferred result. The exact palette matching procedure is up to the artist - our focus in this work is to show that our model respects the palette that was asked for, even extreme palette transfers. We show results for two color-matching algorithms. We call matching based on the most similar color as the $\mathtt{color}$ method; by negating the pairwise distance matrix between source and target colors, we can effectively match based on most dissimilar color, which results in interesting effects ($\mathtt{negative-color}$).
@@ 4 Results & Discussion
Dequantization with Texture: In Fig. 4, we evaluate our models when texture conditioning is available. (top) We are quantizing the GT image and attempting to reconstruct it from its palette, along with additional texture conditioning (i.e. thresholded gradient, gradient, luminance). (bottom) In the case of palette transfers, we do not know the GT image; thus to evaluate, we can quantize the generated image and compare it (i.e. calculate the SSIM/PSNR) with the input (which is a quantized image). Observe how the quality improves with more colors in the palette. Thus, the Palette-SSIM metric is a reasonable proxy when GT is not available, but it’s not perfect. Notice that $\mathsf{ours-T}$ (thresholded gradient) performs better than $\mathsf{ours-L-post}$ (which conditions on GT luminance and enforces it as post-process). Palette error has curious difficulties that follow from the behavior of the quantization algorithm. If we impose the original image’s luminance, we might change the quantization substantially. In Fig. 6 we evaluate our four methods when texture accompanies the quantized image as conditioning. The row indicates the number of colors in the palette; the first column is the quantized image. The subsequent 3 columns show conditioning texture for the 3 models. $\mathsf{ours-L-post}$ and $\mathsf{ours-L}$ condition on luminance, with the former enforcing luminance as a post-process. $\mathsf{ours-G}$ conditions on image gradients and $\mathsf{ours-T}$ conditions on thresholded gradients. All models - even those with less information available - are capable of dequantizing the input quite well, though the luminance-based models tend to achieve better color reproduction.
@@
Table 1: Palette transfer metrics for $\mathtt{color}$ method. Even though $\mathsf{ours-T}$ performed slightly worse in texture-conditioned dequantization (Fig. 4), it excels in palette transfers because thresholded gradients decorrelate local texture from local luminance.
@@ 6 Additional Patch Editing Examples
Quantitative evaluation We select 1000 random test images at 256-res, and pick random patches in each to remove. On average, $13.4\%$ of pixels are masked out. In Table 3, we replace the removed patch with the mean color of its support, and measure how well the model can reconstruct the image. As expected, more texture information is helpful, as shown by the strong numbers for $\mathsf{ours-L}$. However, evaluating each model’s ability to support extreme color changes tells a different story. In Table 4, we repeat this experiment, but we replace each patch with a random $(R,G,B)$ triple. Here, conditioning on texture resulted in the thresholded gradient model performing best, while luminance-conditioning performed worst. This is because conditioning on luminance prevented the network from generating colors with a different luminance that what was asked for in the patch conditioning. Note that in the absence of GT, we use palette error as a proxy error metric. This metric is not perfect, as for example the models without texture conditioning performed slightly better than with texture. However, the palette error gives a strong indication as to whether the model generated an image whose color palette matches what was requested. We use 16 colors in the palette when running this error metric.
@@
$\mathsf{ours-T}$ 30.68 0.9336 29.3 0.921
$\mathsf{ours-G}$ 38.77 0.9823 30.95 0.9315
$\mathsf{ours-L}$ 40.12 0.9845 29.68 0.9202
$\mathsf{ours-L-post}$ 39.82 0.9867 38.27 0.9834
@@
$\mathsf{ours-T}$ 23.56 0.8241 23.87 0.8285
$\mathsf{ours-G}$ 22.37 0.7964 23.58 0.8208
$\mathsf{ours-L}$ 21.76 0.7895 24.07 0.8347
$\mathsf{ours-L-post}$ 20.57 0.7739 20.53 0.7731